# Xu Shen  and  Yuqiang Zheng Morningside Center of Mathematics Academy of Mathematics and Systems Science Chinese Academy of Sciences

### v2

@@ A.3. EKOR stratification
$\operatorname{\mathscr{A}}_{0}\to[\operatorname{\mathcal{G}}_{0}\backslash M^{\mathrm{loc}}].$
@@
Consider $B=L$ a totally real field over $\mathbb{Q}$ and $V=L^{2}$, then $G$ is a subgroup of $\operatorname{Res}_{L|\mathbb{Q}}\operatorname{GL}_{2}$. In this case, the canonical model $\operatorname{\mathscr{A}}$ (with maximal parahoric level at $p$) was constructed in [11] and coincides with $\operatorname{\mathscr{A}}^{\operatorname{naive}}$. Let $x=(A,\lambda,\iota,\alpha)\in\operatorname{\mathscr{A}}_{0}(k)$ and $\operatorname{\mathcal{H}}$, $\omega$ corresponding sheaves as above. For each $i,j$, $\operatorname{\mathcal{H}}_{i,j}\simeq k[\varepsilon_{i}]^{2}$ and $\omega_{i,j}\subset\operatorname{\mathcal{H}}_{i,j}$ is a $k[\varepsilon_{i}]$-submodule with $k$-dimension $e_{i}$. So there is an integer $a_{i,j}$ such that
@@
Now consider the tuple $a=(a_{i,j})$ such that $x\in\operatorname{\mathscr{A}}_{0}^{a}$, then $a_{i,j}$ is the maximal integer such that
@@
If we write $M_{i,j}=\operatorname{\mathcal{H}}_{i,j}/\varepsilon_{i}\operatorname{\mathcal{H}}_{i,j}$, the Ver map $\operatorname{\mathcal{H}}_{i,j+1}\to\operatorname{\mathcal{H}}_{i,j}$ induces a morphism
@@
The type of the $F$-zip $(M,\operatorname{Ker}(V),\operatorname{Ker}(F),\varphi_{\bullet})$ is determined by the dimension of $\operatorname{Ker}(V)$, which is equal to the number of the index $(i,j)$ such that $a_{i,j}\neq e_{i}/2$. More explicitly, consider the function
@@ A.4. Global construction of EKOR stratification
$\operatorname{Sht}_{\mu,K}^{\mathrm{loc}}\to\operatorname{Sht}_{\mu,K}^{\mathrm{loc}(\infty,1)}\to\operatorname{Sht}_{\mu,K}^{\mathrm{loc}(m,1)}\to[\operatorname{\mathcal{G}}_{0}\backslash M^{\mathrm{loc}}].$
@@
$\lambda_{K}:\operatorname{\mathscr{A}}_{0}^{\operatorname{pf}}\to[\operatorname{\mathcal{G}}_{0}\backslash M^{\mathrm{loc}}],$
which is perfectly smooth. The same proof of [57]*Theorem 4.4.3 (there is a small gap in the last step of the proof concerning the involved diagram, which is inherited from the corresponding place of the work of Xiao-Zhu; but a small modification without the commutativity of that diagram will make the argument still work) gives
@@
[…]
@@ A.5. Non-emptiness of EKOR strata
Let $\tau_{\mu}$ be the minimal element in $\operatorname{Adm}(\mu)$ with respect to the Bruhat order. Let $\breve{I}$ be a fixed Iwahori subgroup of $\breve{G}:=G(\breve{\operatorname{\mathbb{Q}}}_{p})$. For simplicity, we may assume that the lattice $\operatorname{\mathscr{L}}$ is determined by a finite index set $J=\{0,\dots,m\}$.
@@
There is a unique basic element $[b_{0}]\in B(G,\mu)$. By [27]*Theorem 1, the basic locus of $\operatorname{\mathscr{A}}_{0}$ is nonempty. Let $x=(A_{j},\lambda_{j},\alpha)_{j\in J}\in\operatorname{\mathscr{A}}(k)$ be a point in the basic locus. Let $D_{j}$ be the Dieudonné module of $A_{j}[p^{\infty}]$ and $N$ the common rational Dieudonné module. Then $D_{j}$ form a lattice chain inside $N$. The Frobenius gives $\delta\in\breve{G}$ such that $[\delta]=[b_{0}]\in B(G,\mu)$.
For any $w\in W_{K}\backslash\widetilde{W}/W_{K}$ and $b\in\breve{G}$, The affine Deligne-Lusztig variety is defined as
@@
by definition. As $\tau_{\mu}$ is a $\sigma$-straight element in $\widetilde{W}$, $\breve{I}\tau_{\mu}\breve{I}$ lies in a single $\sigma$-conjugacy class of $\breve{G}$ by [22]*Theorem 5.1.(a). Moreover, such conjugacy class is given by $[\delta]$. This shows the non-emptiness of $X_{\tau_{\mu}}(\delta)$. ∎
@@
Thus $gD_{j}$ corresponds to a $p$-divisible group which is isogenous to $A_{j}[p^{\infty}]$, so we get an abelian variety $gA_{j}$. The polarization $\lambda_{j}$ and $\operatorname{\mathcal{O}}_{B}$-action extends naturally to $gA_{j}$. The prime to $p$ level structure $\alpha$ also extends to $gA_{j}$, we thus obtain a triple $(gA_{j},g\lambda_{j},\alpha)$. This triple gives a $k$-point $gx\in\operatorname{\mathscr{A}}_{I}(k)$. By construction, $gx$ must live in the minimal EKOR stratum. Then by closure relation, we get the non-emptiness of EKOR strata. ∎

### v3

@@ A.3. EKOR stratification
[…]
@@
[…]
@@
Now consider the tuple $a=(a_{i,j})$ such that $x\in\operatorname{\mathscr{A}}_{0}^{a}(k)$, then $a_{i,j}$ is the maximal integer such that
@@
If we write $M_{i,j}=\operatorname{\mathcal{H}}_{i,j}/\varepsilon_{i}\operatorname{\mathcal{H}}_{i,j}$, the Verschiebung map $\operatorname{\mathcal{H}}_{i,j+1}\to\operatorname{\mathcal{H}}_{i,j}$ induces a morphism
@@
The type of the $F$-zip $(M,\operatorname{Ker}(V),\operatorname{Ker}(F),\varphi_{\bullet})$ is determined by the dimension of $\operatorname{Ker}(V)$, which is equal to the number of the indices $(i,j)$ such that $a_{i,j}\neq e_{i}/2$. More explicitly, consider the function
@@ A.4. Global construction of EKOR stratification
[…]
@@
[…]
which is perfectly smooth. The same proof as [59]*Theorem 4.4.3 (there is a small gap in the last step of the proof concerning the involved diagram, which is inherited from the corresponding place of the work of Xiao-Zhu; but a small modification without the commutativity of that diagram will make the argument still work, cf. [60].) shows
@@
${\operatorname{\mathscr{A}}_{0}^{\operatorname{pf}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{v_{K}}$$\scriptstyle{\lambda_{K}}$${\operatorname{Sht}_{\mu,K}^{\mathrm{loc}(m,1)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{K}}$${{[\operatorname{\mathcal{G}}_{0}\backslash M^{\mathrm{loc}}]}}$
@@ A.5. Non-emptiness of EKOR strata
[…]
@@
There is a unique basic element $b_{0}]\in B(G,\mu)$. By [29]*Theorem 1, the basic locus of $\operatorname{\mathscr{A}}_{0}$ is non-empty. Let $x=(A_{j},\lambda_{j},\alpha)_{j\in J}\in\operatorname{\mathscr{A}}(k)$ be a point in the basic locus. Let $D_{j}$ be the Dieudonné module of $A_{j}[p^{\infty}]$ and $N$ the common rational Dieudonné module. Then $D_{j}$ form a lattice chain inside $N$. The Frobenius gives $\delta\in\breve{G}$ such that $\delta]=[b_{0}]\in B(G,\mu)$.
[…]
@@
by definition. As $\tau_{\mu}$ is a $\sigma$-straight element in $\widetilde{W}$, $\breve{I}\tau_{\mu}\breve{I}$ lies in a single $\sigma$-conjugacy class of $\breve{G}$ by [25]*Theorem 5.1.(a). Moreover, such conjugacy class is given by $\delta]$. This shows the non-emptiness of $X_{\tau_{\mu}}(\delta)$. ∎
@@
Thus $gD_{j}$ corresponds to a $p$-divisible group which is isogenous to $A_{j}[p^{\infty}]$, so we get an abelian variety $gA_{j}$. The polarization $\lambda_{j}$ and $\operatorname{\mathcal{O}}_{B}$-action extends naturally to $gA_{j}$. The prime to $p$ level structure $\alpha$ also extends to $gA_{j}$; we thus obtain a triple $(gA_{j},g\lambda_{j},\alpha)$. This triple gives a $k$-point $gx\in\operatorname{\mathscr{A}}_{I}(k)$. By construction, $gx$ must live in the minimal EKOR stratum. Then by closure relation, we get the non-emptiness of EKOR strata. ∎
Conflicts of interest: none.
Financial Support: the first author was partially supported by the National Key R&\&D Program of China 2020YFA0712600, the CAS Project for Young Scientists in Basic Research, Grant No. YSBR-033, and the NSFC grant No. 12288201.